OpenStudy (anonymous):

what is the standard form of Ellipse with foci at (4, 2) and (10, 2) and co-vertices at (7, 0) and (7, 4)

OpenStudy (anonymous):

The numerator and denominator of a fraction are in the ratio of 3 to 5. If the numerator and denominator are both increased by 2, the fraction is now equal to 2/3. If n = the numerator and d = the denominator, which of the following systems of equations could be used to solve the problem? 5n = 3d and 3n + 6 = 2d + 4 5n = 3d and 4n + 4 = 3d + 6 3n = 5d and 3n + 6 = 2d + 4

OpenStudy (anonymous):

JOHN SHUT UP

OpenStudy (anonymous):

an how do uknow my name

OpenStudy (anonymous):

ok honey so the center is 7,2 becuz it is the midpoint of the foci. now we kno that its ellipse so c^2 = a^2 - b^2 so just plug and chug baby. so c we kno is 3 and b is 2 so we get 9 = a^2 - 4, then we get a = plug or minus sqrt of 13. so the final answer is :